OpenStudy (anonymous):

how do you find the absolute value of |5+12i|

jimthompson5910 (jim_thompson5910):

Hint: |a+bi| = sqrt(a^2 + b^2)

OpenStudy (anonymous):

the answer is 13=D

OpenStudy (anonymous):

how do i get 13 ?

OpenStudy (anonymous):

do what jim told you to do fallow the formula

OpenStudy (anonymous):

not good with math though so i dont know

OpenStudy (anonymous):

\(\sqrt{5^2+12^2}\) a familiar right triangle

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):